Train A is traveling at 60 mph and train B is traveling 80 mph. Train A passes a station at 9:10 pm. If train B passes the same station at 9:25 pm, at that time will train B catch up to train A?
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

60T=80(T-.25)
60T=80T-20
60T-80T=-20
20T=20
T=20/20
T=1 HOUR AFTER TRAIN A LEAVES THEY WILL MEET.
9:10+1:00=10:10 TIME OF THEIR MEETING.
PROOF:
60*1=80(1-.25)
60=80*.75
60=60

Answer by Alan3354(69443)

Do these the easy way:
In 15 minutes, train A goes 15 miles (60 mph x 1/4 hour)
Train B gains on train A at 20 mph (80 - 60)
15 miles/20 mph = 3/4 hour = 45 minutes
9:25 + 45 = 10:10